# A Disproof of Large Language Model Consciousness: The Necessity of Continual Learning for Consciousness

### 12 Jan 2026 v2

@@ 3 A Formal Framework for Falsifying Theories of Consciousness
A significant amount of confusion about the Kleiner-Hoel dilemma was simply not understanding that the dilemma has this second horn of strict dependency. E.g., in one of the only cases in the literature of attempting to confront the problem head-on, the criticism ended up being that the substitution argument could not possibly be true, because we did not factor in needing to change the predictions for functionalist theories of consciousness after substitutions [34]. But in arguing that predictions must change just the same way as inferences following a substitution, the paper proposes a trivial theory of consciousness (a kind of global input/output functionalism) that falls on the other horn of the Kleiner-Hoel dilemma, thus proving our point that the space is indeed hard to navigate.
In what follows, the Kleiner-Hoel formalism for falsification of theories of consciousness is used to show several results. In the next section, I show that it provides a disproof of LLM consciousness. After that, I present a case of avoiding the dilemma in humans but not LLMs, indicating that the actual conclusion of the Kleiner-Hoel dilemma is that it disproves consciousness in LLMs but not in humans.
@@ 4 The Proximity Argument
Here, I assume that trivial theories of consciousness must be false (rather than just trivial), since otherwise there is no scientifically informative theory of consciousness. This assumption makes the language and conclusions easier to follow throughout (e.g., instead of writing “a system for which there are no non-trivial theories of consciousness available” I can say “a non-conscious system”); however, all the conclusions would still hold but with this sort of replacement language if this assumption (which is just that a scientifically-informative theory of consciousness actually exists) were not true.
@@
A lookup table offers an archetypal example of a non-conscious system under Definition 4.2. Any lookup table $L$ simply implements a input/output function $f:I\rightarrow O$. Imagine a theory $T$ that predicts $L$ has non-trivial consciousness. What could $T$ base its prediction of consciousness on? $L$ is not internally complex, lacks all dynamics, is entirely static, has no memory capabilities, and there is no information flow across any parts or modules. Certainly, a theory cannot make meaningful predictions of consciousness based on merely the fact that some set of IF-THEN statements exist. The sole relevant property for a theory of consciousness pertaining to $L$ is the input/output of the function $f$ itself. However, that’s the same function that determines all the experimental inference data! Therefore, any $T$ predicting $L$ is conscious must base this prediction solely on $f$, which is identical to the basis of inference, creating strict dependency. Thus any given $T$ is trivial. Thus, $L$ is necessarily a non-conscious system (as long as, per definition 4.1, non-trivial theories of consciousness exist).
@@
For a pair of systems, $s$ and $s^{\prime}$, that share identical input/output (or approximately equivalent input/output, wherein this could describe behavior or responding to text, etc.), the substitution distance $d_{S}(s,s^{\prime})$ is determined by the space of properties $\mathcal{P}$ that differ between $s$ and $s^{\prime}$.
As an example of Definition 4.3, consider the well-detailed case of unfolding a recurrent artificial neural network (RNN) into a feedforward artificial neural network (FNN) that implements the same input/output function, as in [31]. In such an unfolding, many properties are lost. E.g., while a single-hidden-layer FNN can approximate any continuous function to arbitrary precision [28, 39], it lacks recurrence, temporal dynamics, higher-order structure, broadcasting, depth, etc. On the other hand, some things remain unchanged, like its substrate and activation functions (i.e., it is still an artificial neural network with individual artificial neurons and connections, running on a computer, etc.). The scope of these lost properties defines the substitution distance (Definition 4.3).
An alternative comparison might be the human brain. Can a human brain be substituted with a functionally-equivalent single-hidden-layer FNN, even one impractically astronomically wide? It is perhaps possible (depending on the notion of possibility)—let us not take sides right now—but the point of this definition is that, if such a substitution were possible, the number of properties “lost” would be much larger (e.g., the brain is biological, contains things like local field potentials, dendritic mechanics, molecular properties, etc.). And the same goes for substituting a human brain with the shortest program capable of carrying out its input/output function (as in Theorem 3.1); much more would be lost than just substituting one computer program for another.
@@
Step 1. Let $M$ be a deployed static LLM, let $N$ be a static single-hidden-layer feedforward neural network implementing the same input/output function as $M$, and let $L$ be a lookup table also implementing the same input/output function.
Step 2. By the universal approximation theorem, any input/output function (like text responses) computed by $M$ can be approximated arbitrarily well by a sufficiently-wide static single-hidden-layer network $N$. By definition, any finite function computed by $N$ can in turn also be implemented as a lookup table, $L$.
Step 3. Therefore there exists a chain of universal substitutions preserving input/output:
@@
At this point in our reasoning, we should already have strong reasons to question LLM consciousness, simply due to how significant this constraint is. Certainly, there are some details like scaffolding, interfaces, if it is deployed, and what type of contemporary LLM, etc., that I am leaving aside. However, at least when it comes to the core “bare bones” of a baseline contemporary LLM (like, say, the original GPT-4, with no router or other complications), any claim to consciousness is confined to its distance to a static single-hidden-layer approximation of its input/output (which is guaranteed by the universal approximation theorem [39]). And note: Here, the substitution of a static single-hidden-layer FNN is another kind of artificial neural network that computes input/output in much the same way, in that it has similar activation functions, is also instantiated on a computer, has identical input/output, and so on. So if we just stopped here at Corollary 4.7, LLM consciousness would, at minimum, be highly constrained. It must require something lost in that chain of universal substitutions (like, say, extra layers).
If one still holds that a single-hidden-layer static FNN network could be non-trivially conscious, then there is also the distance to the provably-trivial lookup table to consider (there, the only relevant property appears to be compression, which is also not a good target for theories of consciousness—following this logic, see Corollary 4.10 for a proof that static single-hidden-layer static FNNs are non-conscious). If a theory of consciousness cannot specify what is important and lost along this substitution chain, then it cannot claim that LLMs are conscious.
@@
Step 1. The substitution chain $L\leftrightarrow N\leftrightarrow M$ is constructed to preserve input/output. Since empirical inferences about consciousness for entities like LLMs are derived from input/output functions (not from, e.g., how many layers they possess), all inference data is fixed across the chain.
@@
Step 4. The only alternative is if $\rho$ does not vary, in which case, $T$ must necessarily ground predictions solely in the input/output structure itself. This creates a situation of strict dependency, rendering $T$ trivial.
@@
An alternative summary of Proposition 4.8 is to say that the substitution chain $L\leftrightarrow N\leftrightarrow M$ is constructed to fix input/output, which essentially exhausts all relevant inference data (especially in light of Corollary 4.10, which proves static single-hidden-layer FNNs are non-conscious), since there would have to be a priori neutral and non-empirical and non-theory-based reasons to assume LLM consciousness, which somehow do not apply to single-hidden-layer FNN consciousness, and to such a degree that it (somehow) overrides the pathological mismatches with a non-conscious system. Therefore, any property that varies across the chain should be independent of inferences. This eliminates the possibility of overcoming the Kleiner-Hoel dilemma within this substitution chain.
@@
By the Constraint Theorem, any non-trivial theory predicting LLMs are conscious must base its predictions on properties in $d_{S}(M,N)\cup d_{S}(N,L)$. By Proposition 4.8, no property in this space is not falsified by the universal substitution chain itself. Therefore, under Definition 4.2, LLMs are non-conscious systems. ∎
@@
Examining a single step in the Proximity Principle: consider the substitution distance $d_{S}(N,L)$ between a lookup table, $L$, and a static single-hidden-layer FNN, $N$, implementing the same input/output function. This distance can be entirely summarized as the degree of compression of the input/output function, and since compression can be varied via universal substitutions, therefore single-hidden-layer FNNs are non-conscious systems.
@@
Step 1. A lookup table with $n$ discrete inputs can be implemented as a maximally wide static single-hidden-layer FNN (with one hidden unit per possible input/output).
Step 2. Substitute in a narrower network with identical input/output. Since that input/output is fixed, the only difference is in the degree of compression in the implementation.
Step 3. Any theory $T$ that grounds its predictions of consciousness in compression can therefore yield independent predictions across networks, all while inferences (necessarily derived from input/output) remain fixed—thus creating the pathological mismatches that falsify $T$.
Step 4. Therefore, only the input/output function itself remains to actually ground $T$, and so by Definition 4.2 of non-conscious systems, static single-hidden-layer FNNs are non-conscious systems. ∎
@@
Zooming out: there exist in the literature previous “no-go” results on AI consciousness, such as arguments against AI consciousness based on causal exclusion [46] (though similar arguments have been given against the efficacy of human consciousness as well [44]), as well as other arguments for biological naturalism [66, 21]. However, one unique aspect of the Proximity Argument is that it makes minimal assumptions other than that: (a) a theory of consciousness cannot be trivial or a priori falsified, (b) that certain universal substitutions are possible (which we know from existent mathematical theorems to be true), and (c) that of widespread or universal mismatches from a theory in the comparative testing setup described would falsify that theory. Nor does it rely on a particular theory; e.g., while transformers provably are entirely feedforward and have zero integrated information (thus, IIT would say that LLMs are not conscious) [5], the Proximity Argument goes beyond relying on a particular theory. It requires no opinion on whether qualia is causally relevant or epiphenomenal, nor whether functionalism is true, nor any other philosophical position about consciousness, nor does it rely on any particular theory of consciousness or even class of theories to be true (it does not, e.g., require solving the Hard Problem [22]). It is a philosophically neutral proof that both constrains and disproves LLM consciousness based on falsification, linking up to historical critiques about functionalism and intelligence (e.g., the Blockhead thought experiment) [16].
@@
It should also be noted that, following Theorem 3.1, aspects of the Proximity Argument could be restated in terms of the shortest program $K(f)$ acting as the substitute for the LLM. Indeed, the shortest program is much like a mirror of a lookup table: instead of being the maximally-complex implementation (wherein every part relates to the function, at least), the shortest program is the minimally-complex implementation, and it too is (a) different from whatever original program or system was specified, and (b) just as much a direct function of input/output.
4.1 The Proximity Argument is Particularly Forceful for LLMs
The Proximity Argument is based on the formal falsification framework developed in [45]; however, while applying it to LLMs, the Proximity Argument does not necessarily require the full acceptance of the Kleiner-Hoel dilemma applying in humans for all possible theories of consciousness. Humans are much less proximate (Proposition 4.5) in “substitution distance” (Definition 4.3) to trivially non-conscious systems, therefore the Constraint Theorem (Theorem 4.6) doesn’t bound the possibility space to the same degree, as there are many more properties lost in universal substitutions. This means there are far more available choices to ground non-trivial theories of consciousness in (although whether that can escape both horns of the dilemma is another matter). At least, with humans, we have good self-introspective reasons to believe we are conscious; theories grounded in phenomenology may even be different from normal scientific theories due to us having first-person evidence [67]. However, grounding theories in phenomenology alone likely requires sacrifices (like falsifiability), and, in light of the following evidence that some theories of consciousness do escape logical issues around falsifiability, this option is set aside here.
A related objection is that we should artificially restrict our scope of consciousness science to apply solely in humans to avoid the Kleiner-Hoel dilemma [3, 72]. This too comes with significant costs, like: (a) substitutions may be harder to a priori identify, but will still apply in humans simply due to independence between inferences and predictions (e.g., [29]), since the difference between a “substitution” and an “advanced experiment” seems impossible to specify; (b) theories lose their universality and this calls into question any extrapolation beyond humans, which represents a large domain for which a science of consciousness should have answers; (c) it is equivalent to saying that, beyond humans, inferences based on general conscious-like behavior contain zero actionable scientific information and cannot be used for inferences about consciousness.
It is (c) that is actually most relevant here, as it makes the Proximity Argument particularly strong: If one tried to get around the Proximity Argument by arguing to solely confirm/falsify theories of consciousness in humans to try to preempt issues of falsification (all while ignoring (a), that substitutions still likely apply), this move would necessarily imply that the intelligent outputs of LLMs contain zero information about their consciousness to begin with and cannot be used for inferences; therefore, we would have no reason to believe LLMs are conscious.
Luckily, these debates can be sidestepped entirely by examples of lenient dependency that avoid the Kleiner-Hoel dilemma in humans but not LLMs. If this can be shown to hold, meaning that we have reasons to believe the Kleiner-Hoel dilemma doesn’t apply to humans under particular classes of theories of consciousness, we have even stronger reasons to think the Proximity Argument rules out consciousness in LLMs, since there can be no argument that it rules out consciousness research altogether.
Therefore, the second part of this paper will explore the requirements for theories to avoid the Kleiner-Hoel dilemma in humans, a property herein referred to as “lenient dependency.”
@@ 4.1 The Proximity Argument is Particularly Forceful for LLMs
First, I argue that lenient dependency must be a continually-existing criterion for theories of consciousness (a high bar). Then, I propose that learning is a form of lenient dependency in humans, and that this is necessarily restricted to truly continual learning (the learning occurring with, or potentially even constituting, every single conscious experience). If this further argument is correct, it would significantly constrain consciousness in humans to involve the process of continual learning. I argue that this restriction is scientifically positive: while it rules out most existing theories in the field, it still points us toward evocative and rich theories of consciousness and reveals what a minimal falsifiable and non-trivial theory of consciousness would look like. Additionally, it indicates that a fundamental difference between LLMs and humans is the ability to continually learn and that LLM’s lack of consciousness may be tied to their static nature.
5 The Necessity of Continual Lenient Dependency

### 19 Jan 2026 v3

@@ 3 A Formal Framework for Falsifying Theories of Consciousness
A significant amount of confusion about the Kleiner-Hoel dilemma was simply missing that the dilemma has this second horn of strict dependency. E.g., in one of the only cases in the literature of attempting to confront the problem head-on, the criticism ended up being that the substitution argument could not possibly be true, because it did not factor in needing to change the predictions for functionalist theories of consciousness after substitutions [42]. But in arguing that predictions must change just the same way as inferences following a substitution, the paper proposes a trivial theory of consciousness (a kind of global I/O functionalism) that falls on the second horn of the Kleiner-Hoel dilemma.
In the next section, I show that the Kleiner-Hoel dilemma provides a particularly strong disproof of LLM consciousness. After that, I present a case of avoiding the dilemma in humans but not LLMs, indicating that it disproves consciousness in LLMs but not in humans.
@@ 4 The Proximity Argument
Here, I assume that trivial theories of consciousness must be false (rather than just trivial), since otherwise there is no scientifically informative theory of consciousness. This assumption makes the language and conclusions easier to follow throughout (e.g., instead of writing “a system for which there are no non-trivial theories of consciousness available” I can say “a non-conscious system”); however, all the conclusions would still hold but with this sort of replacement language if this assumption (which is just that a scientifically-informative theory of consciousness actually exists) were not true. See subsection 4.1 for relevant discussion about rejecting this and instead accepting trivial theories as true.
@@
A lookup table offers an archetypal example of a non-conscious system under Definition 4.2. Any lookup table $L$ simply implements an I/O function $f:I\rightarrow O$. Imagine a theory $T$ that predicts $L$ has non-trivial consciousness. What could $T$ base its prediction of consciousness on? $L$ is not internally complex, lacks all dynamics, is entirely static, has no memory capabilities, and there is no information flow across any parts or modules. Certainly, a theory cannot make meaningful predictions of consciousness based on merely the fact that some set of IF-THEN statements exist. The sole relevant property for a theory of consciousness pertaining to $L$ is the I/O of the function $f$ itself. However, that’s the same function that determines all the experimental inference data! Therefore, any $T$ predicting $L$ is conscious must base this prediction solely on $f$, which is identical to the basis of inference, creating strict dependency. Thus any given $T$ is trivial. Thus, $L$ is necessarily a non-conscious system (as long as, per Definition 4.1, non-trivial theories of consciousness exist).
@@
For a pair of systems, $s$ and $s^{\prime}$, that share identical I/O (or approximately equivalent I/O, wherein this could describe behavior or responding to text, etc.), the substitution distance $d_{S}(s,s^{\prime})$ is determined by the space of properties $\mathcal{P}$ that differ between $s$ and $s^{\prime}$.
As an example of Definition 4.3, consider the well-detailed case of unfolding a recurrent artificial neural network (RNN) into a feedforward artificial neural network (FNN) that implements the same I/O function, as in [37]. In such an unfolding, many properties are lost. E.g., while a single-hidden-layer FNN can approximate any continuous function to arbitrary precision [32, 48], it lacks recurrence, temporal dynamics, higher-order structure, broadcasting, depth, etc. On the other hand, some things remain unchanged, like its substrate and activation functions (i.e., it is still an artificial neural network with individual artificial neurons and connections, running on a computer, etc.). The scope of these lost properties defines the substitution distance (Definition 4.3).
@@
Step 1. Let $M$ be a deployed static LLM, let $N$ be a static single-hidden-layer feedforward neural network implementing the same I/O function as $M$, and let $L$ be a lookup table also implementing the same I/O function.
Step 2. By the universal approximation theorem, any I/O function (like text responses) computed by $M$ can be approximated arbitrarily well by a sufficiently-wide static single-hidden-layer network $N$. By definition, any finite function computed by $N$ can in turn also be implemented as a lookup table, $L$.
Step 3. Therefore there exists a chain of universal substitutions preserving I/O:
@@
At this point in our reasoning, we should already have strong reasons to question LLM consciousness, simply due to how significant this constraint is. Certainly, there are some details like scaffolding, interfaces, if it is deployed, and what type of contemporary LLM, etc., that I am leaving aside. However, at least when it comes to the core “bare bones” of a baseline contemporary LLM (like, say, the original GPT-4, with no router or other complications), any claim to consciousness is confined to its distance to a static single-hidden-layer approximation of its I/O (which is guaranteed by the universal approximation theorem [48]). And note: Here, the substitution of a static single-hidden-layer FNN is another kind of artificial neural network that computes I/O in much the same way, in that it has similar activation functions, is also instantiated on a computer, has identical I/O, and so on. So if we just stopped here at Corollary 4.7, LLM consciousness would, at minimum, be highly constrained. It must require something lost in that chain of universal substitutions (like, say, extra layers).
If one still holds that a single-hidden-layer static FNN is non-trivially conscious, then there is also the distance to the provably-trivial lookup table to consider (there, the only relevant property appears to be compression, which is also not a good target for theories of consciousness—see Corollary 4.10 for a proof that static single-hidden-layer static FNNs are non-conscious). If a theory of consciousness cannot specify what is important and lost along this substitution chain, then it cannot claim that LLMs are conscious.
@@
Step 1. The substitution chain $L\leftrightarrow N\leftrightarrow M$ is constructed to preserve I/O. Since empirical inferences about consciousness for entities like LLMs are derived from I/O functions (not from, e.g., how many layers they possess), all inference data is fixed across the chain.
@@
Step 4. The only alternative is if $\rho$ does not vary, in which case, $T$ must necessarily ground predictions solely in the I/O structure itself. This creates a situation of strict dependency, rendering $T$ trivial.
@@
An alternative summary of Proposition 4.8 is to say that the substitution chain $L\leftrightarrow N\leftrightarrow M$ is constructed to fix I/O, which essentially exhausts all relevant inference data (especially in light of Corollary 4.10, which proves static single-hidden-layer FNNs are non-conscious), since there would have to be a priori neutral and non-empirical and non-theory-based reasons to assume LLM consciousness, which somehow do not apply to single-hidden-layer FNN consciousness, and to such a degree that it (somehow) overrides the pathological mismatches with a non-conscious system. Therefore, any property that varies across the chain should be independent of inferences. This eliminates the possibility of overcoming the Kleiner-Hoel dilemma within this substitution chain.
@@
By the Constraint Theorem, any non-trivial theory predicting LLMs are conscious must base its predictions on properties in $d_{S}(M,N)\cup d_{S}(N,L)$. By Proposition 4.8, every property in this space is falsified by the universal substitution chain itself. Therefore, under Definition 4.2, LLMs are non-conscious systems. ∎
@@
Examining a single step in the Proximity Principle: consider the substitution distance $d_{S}(N,L)$ between a lookup table, $L$, and a static single-hidden-layer FNN, $N$, implementing the same I/O function. This distance can be entirely summarized as the degree of compression of the I/O function, and since compression can be varied via universal substitutions, therefore single-hidden-layer FNNs are non-conscious systems.
@@
Step 1. A lookup table with $n$ discrete inputs can be implemented as a maximally wide static single-hidden-layer FNN (with one hidden unit per possible I/O).
Step 2. Substitute in a narrower network with identical I/O. Since that I/O is fixed, the only difference is in the degree of compression in the implementation.
Step 3. Any theory $T$ that grounds its predictions of consciousness in compression can therefore yield independent predictions across networks, all while inferences (necessarily derived from I/O) remain fixed—thus creating the pathological mismatches that falsify $T$.
Step 4. Therefore, only the I/O function itself remains to actually ground $T$, and so by Definition 4.2 of non-conscious systems, static single-hidden-layer FNNs are non-conscious systems. ∎
@@
Zooming out: there exist in the literature previous “no-go” results on AI consciousness, such as arguments against AI consciousness based on causal exclusion [58] (though similar arguments have been given against the efficacy of human consciousness as well [56]), as well as other arguments for biological naturalism [85, 21]. However, one unique aspect of the Proximity Argument is that it makes minimal assumptions other than that: (a) a theory of consciousness cannot be trivial or a priori falsified, (b) that certain universal substitutions are possible (which we know from existent mathematical theorems to be true), and (c) that of widespread or universal mismatches from a theory in the comparative testing setup described would falsify that theory. Nor does it rely on a particular theory; e.g., while transformers provably are feedforward and have zero integrated information (thus, IIT would say that LLMs are not conscious) [5], the Proximity Argument goes beyond relying on a particular theory. It requires no opinion on whether qualia is causally relevant or epiphenomenal, nor whether functionalism is true, nor any other philosophical position about consciousness, nor does it rely on any particular theory of consciousness or even class of theories to be true (it does not, e.g., require solving the Hard Problem [24]). It is a philosophically neutral proof that both constrains and disproves LLM consciousness based on falsification, linking up to historical critiques about functionalism and intelligence (e.g., the Blockhead thought experiment [16] and the Chinese Room thought experiment [81]) although it is not framed as a reductio ad absurdum but based on concerns about falsifiability.
@@
It should also be noted that, following Theorem 3.1, aspects of the Proximity Argument could be restated in terms of the shortest program $K(f)$ acting as the substitute for the LLM. Indeed, the shortest program is much like a mirror of a lookup table: instead of being the maximally-complex implementation (wherein every part relates to the function, at least), the shortest program is the minimally-complex implementation, and it too is (a) different from whatever original program or system was specified, and (b) just as much a direct function of I/O.
4.1 The Kleiner-Hoel Dilemma is Particularly Forceful for LLMs
Why does the Kleiner-Hoel dilemma not apply just as much to humans? Here, I give multiple reasons why the arguments are particularly forceful in LLMs while considering possible interpretations of these proofs. I also consider cases where the Kleiner-Hoel dilemma does apply in humans, how theories might escape the dilemma, and how rescuing LLM consciousness requires extreme and unscientific beliefs about consciousness.
  1. 1.
    Non-conscious universal substitutions for LLMs are well-defined and constructible. In some cases, universal substitutions may be merely conceivable (e.g., consider an infinite lookup table implementing $f$, wherein $f$ is somehow given). However, in practice, this apparent conceivability may turn out to be ill-defined. E.g., can a human brain be substituted with a functionally-equivalent single-hidden-layer FNN, even one impractically astronomically wide? Is a human’s f clearly definable? Asking these questions explicitly is not necessarily agreeing that universal substitutions are ill-defined in humans—indeed, the entire rest of the paper will assume these questions are answerable. However, it is undeniable there is significant subtlety and complexity to them. It may be that actual definable and constructible universal substitutions for humans (especially those that count as non-conscious according to Definition 4.2) could be somehow ruled out by details we don’t know or are hard to foresee (e.g., quantum processes in the brain [44], or some computational irreducibility in the process [92], or the no-cloning theorem [93] or some other factor that crops up in an unexpected way).
    However, for baseline LLMs, it appears from the literature that universal substitutions are actually constructible via chaining together known methods. Specifically, it is known that transformers can be arbitrarily approximated as RNNs [55, 27]. Then, it is known that (static) RNNs can be “unfolded” into a single-hidden-layer FNN [37, 75], which can in turn be replaced via enumeration via a lookup table. So the chain discussed in the Proximity Argument (e.g., Theorem 4.6) is literally constructible via known methods, and could potentially even be given as defined transformations. Perhaps for a small-enough language model, some version of it could be transformed this way with current technology. That is, we seem to have a general understanding of how to implement non-conscious (Definition 4.2) universal substitutions for LLMs. This should call into question their consciousness significantly.
  2. 2.
    The potential for theories that evade the Kleiner-Hoel dilemma in humans. Due to the Proximity Principle (Proposition 4.5) there is a far greater Substitution Distance (Definition 4.3) between humans and available non-conscious substitutions (which, from now on, I assume are well-defined and possible). This means that humans have more Consciousness-Relevant Properties (Definition 4.4). Meanwhile, LLMs are much more constrained via the Constraint Theorem (Theorem 4.6). This means there are far more options for theories of consciousness in humans compared to LLMs. Importantly, this means there are far more potential theories that demonstrate “lenient dependency” in that they avoid the Kleiner-Hoel dilemma but still are falsifiable and contain meaningful scientific information.
@@ 4.1 The Kleiner-Hoel Dilemma is Particularly Forceful for LLMs
    The remaining sections of this paper will deal with exploring the nature of this lenient dependency via a class of theories of consciousness that are sensitive to or grounded in properties of continual learning.
  3. 3.
    The greater potential for theories that are unscientific or unfalsifiable to assign consciousness to humans. The alternative to rejecting trivial theories and thus denying conclusions like Definition 4.2 of Non-conscious Systems, is to accept trivial theories (and only trivial theories; otherwise, they would be falsified by the first horn of the Kleiner-Hoel dilemma). And there are indeed various theories of consciousness that may contain minimal scientific information and so avoid the Kleiner-Hoel dilemma by being trivial. These might be categorized as metaphysical theories that are also scientifically trivial, as per Definition 4.1 of trivial theories. However, instead of rejecting such theories in favor of a scientifically testable one, as here, one might imagine accepting them. Even so, all such theories have far more “room” in humans, due to the many more properties of humans that can serve as Consciousness-Relevant Properties (Definition 4.4). It is beyond the scope of this paper to delineate metaphysical theories that are still scientifically trivial per Definition 4.1; however, potential candidates include theories like substance dualism [35] or analytic idealism [54]. Even panpsychist theories like Russellian monism [23] face a combination problem [73] that seems more solvable in an integrated and plastic human brain than a feedforward and static LLM. So even when advocating for a trivial theory of consciousness, belief in LLM consciousness is constrained to a subset of theories which seem the least metaphysically “worthwhile” (e.g., behaviorism).
  4. 4.
    The scope of inferences could be restricted. One potential strategy for dealing with the Kleiner-Hoel dilemma is to severely restrict the scope of consciousness science. It has been advocated to simply test theories in humans and ignore contradictory reports from substitutions [3, 90]. It is imagineable that this could be justifiable if consciousness science is unlike other sciences and relies on first-person evidence [86]. In particular, this fits with theories based in phenomenology, like IIT [77]. However, it is already known that such theories show that LLMs would possess zero integrated information [5], and IIT would rule out even a simulated brain on a computer being conscious [38]. Overall, while this seems conceptually possible, this move is restrictive. A theory of consciousness would be not universal or robust, but highly fragile to falsification, and kept from falsification via artificial restrictions. Currently, it also remains undefined. What, ultimately, is the difference between an experiment and a substitution? Already there have been proposed experiments in humans that complexify any attempted demarcation [34]. And most relevantly to the thesis here, restricting inferences is particularly problematic if someone were to advocate for it and also LLM consciousness at the same time, as it means accepting that non-human behavior (even highly complex) contains zero actionable scientific information about consciousness. Therefore, there would be no particular reason to think LLMs would possess consciousness to begin with, and no behavior from an LLM would count as evidence for its consciousness.
  5. 5.
    The Kleiner-Hoel dilemma could indicate the undecidability between theories of consciousness. I have previously argued that consciousness could act much like a Gödel sentence in science [47]—there are alternative proposals for various forms of related mysterianism as well [72]. This would likely mean there are no falsifiable theories of consciousness, and, if the Kleiner-Hoel dilemma was truly universal across systems (including humans) and theories of consciousness (and only if this universality held) it would support that consciousness is beyond empirical science’s ken; however, the reasoning for metaphysical but unscientific trivial theories having more “room” for humans in them would still apply (as in #2).
@@
For the rest of the paper, I explore the scientifically-promising initial given reason for why the Kleiner-Hoel dilemma may not apply in humans: theories of consciousness may possess lenient dependency. In the following section, I argue that lenient dependency must be a continually-existing criterion for theories of consciousness (a high bar). Then, I propose that learning is a form of lenient dependency in humans, and that this is necessarily restricted to truly continual learning (the learning occurring with, or potentially even constituting, every single conscious experience). If this further argument is correct, it would significantly constrain consciousness in humans to involve the process of continual learning. I argue that this restriction is scientifically positive: while it rules out most existing theories in the field, it still points us toward evocative and rich theories of consciousness and reveals what a minimal falsifiable and non-trivial theory of consciousness would look like. Additionally, it indicates that a fundamental difference between LLMs and humans is the ability to continually learn and that LLM’s lack of consciousness may be tied to their static nature.
5 Continual Lenient Dependency